OpenStudy (anonymous):

find dy/dx by implicit differentiation (1+sin^3(xy^2))^1/2= y

OpenStudy (schrodingers_cat):

Well all I have to say is chain rule So, Start with u = 1+sin^3(xy^2) then d/du( sqrt(u)) = 1/(2(sqrt(u))) then take derivative with respect to x of 1+sin^3(xy^2) = 3sin(xy^2)^2 At this point you'll have , (3sin(xy^2)^2)(d/dx(sin(xy^2))/2(sqrt((1+sin^3(xy^2))) = y' then take derivative with respect to (d/dx(sin(xy^2) = (cos(xy^2)) Then take derivative of interior and you will get d/dx(xy^2) = y^2 + 2xyy' So, you will have this in total (3sin(xy^2)^2)(cos(xy^2))(y^2 + 2xyy')/2(sqrt((1+sin^3(xy^2))) = y' I will leave it to you to solve for y' because at this point it is just algebra :)